# Global Compression Commander: Plug-and-Play Inference Acceleration for High-Resolution Large Vision-Language Models

### v4

@@ 4.1 Adaptive Compression Adjustment
As presented in the bottom-left of Figure 4, GlobalCom2 first analyzes the semantic contribution of each local crop to the overall visual information, and accordingly applies adaptive token compression at different levels.
To achieve differentiated token compression for local crops, GlobalCom2 adaptively adjusts the compression degrees for each crop based on its informativeness. Specifically, for the $j$-th crop, we define its information richness score $s_{j}^{G}$ from the global perspective by accumulating patch-to-[CLS] attention scores over its corresponding region in the global thumbnail: $s_{j}^{G}=\sum_{i\in\text{crop}_{j}}s^{G}_{i}$, where $s^{G}_{i}$ denotes the informativeness score of each token in the crop region from the thumbnail. To stabilize $s_{j}^{G}$, we compute $\tilde{s}_{j}=(s_{j}^{G}-\max(s_{j}^{G}))/\tau$ with a temperature parameter $\tau=10$, and then compute the relative importance weight $\sigma_{j}$ of the $j$-th crop via softmax across all $n$ crops:
@@
where $\epsilon=10^{-8}$ prevents division by zero. Finally, we adjust the preset retention ratio $R(\%)$ according to the computed importance weight $\sigma_{j}$. This adjustment ensures crops with higher importance preserve more tokens while maintaining the average retention ratio:
@@
where $\sigma_{j}-\frac{1}{n}$ represents the relative deviation from average importance, leading to more aggressive compression for less important regions (values exceeding 100% are capped).
To this end, GlobalCom2 optimally allocates compression degrees (i.e., $\{r_{1},r_{2},r_{3},r_{4}\}$ in Figure 4) based on each crop’s information richness from global view, enabling differentiated token compression degrees across crops.
@@ 4.2 Holistic Token Evaluation
After determining the optimal compression budgets, GlobalCom2 evaluates the importance of tokens within each crop and preserves the essential ones, as shown in the bottom-right of Figure 4.
After independent ViT encoding of each local crop, the attention values between patch tokens and the [CLS] token in its corresponding attention map reflect each token’s semantic importance within the crop. Similar to global analysis, GlobalCom2 computes a set of local importance scores $\{s_{j}^{L}\}_{j=1}^{n}$, where $s_{j,i}^{L}$ denotes the score of the $i$-th token in the $j$-th crop based on its own final-layer attention map. Here, each $s_{j}^{L}$ represents the attention scores after excluding padding tokens. However, since local crops are encoded independently during vision encoding stage, these local scores only reflect token importance within individual crops, failing to capture their significance in the full image context. This local-only evaluation poses a potential challenge, as semantically crucial elements that bridge different image regions might be overlooked.
Motivated by observation that global thumbnails provide holistic visual representations, GlobalCom2 extends local token evaluation with global context to achieve more comprehensive importance assessment. As shown in Figure 4, GlobalCom2 reshapes the 1D attention scores ${s}^{G}$ of $\mathbf{X}^{G}$ into 2D format and bilinearly interpolates them to match the original image dimensions (denoted as “interp.”), yielding $\hat{s}^{G}\in\mathbb{R}^{H\times W}$. Subsequently, $\hat{s}^{G}$ is split into $n$ sub-attention maps $\{\hat{s}_{j}^{G}\}_{j=1}^{n}$ according to the local crop partitioning, where these sub-attention maps measure the semantic significance of tokens in local crops from a global perspective. Combining sub-attention maps $\{\hat{s}_{j}^{G}\}_{j=1}^{n}$ with the local importance scores $\{s_{j}^{L}\}_{j=1}^{n}$, GlobalCom2 is able to comprehensively evaluate token importance from both global and local perspectives.
Specifically, both global sub-attention maps $\{\hat{s}_{j}^{G}\}_{j=1}^{n}$ and local attention scores $\{s_{j}^{L}\}_{j=1}^{n}$ are first L2-normalized. The holistic score $s_{j,i}$ for the $i$-th token in the $j$-th local crop is then calculated by integrating its normalized global importance score $\hat{s}_{j,i}^{G}$ and local importance score $s_{j,i}^{L}$ as:
$s_{j,i}=\alpha\hat{s}_{j,i}^{G}+(1-\alpha)s_{j,i}^{L}$ (5)
Here, $\alpha\in[0,1]$ balances the two importance scores, defaulting to 0.5 for equal weighting. Based on the holistic scores, the token compression process for each local crop can be formulated as:
$\mathbf{X}^{L}_{j}\rightarrow\mathbf{\hat{X}}^{L}_{j}=\text{TopK}(\mathbf{X}^{L}_{j},s_{j},r_{j}\times N)$ (6)
where $\{\mathbf{\hat{X}}^{L}_{j}\}_{j=1}^{n}$ represents the compressed token sequences. Through holistic token evaluation, GlobalCom2 identifies tokens that are significant in global context while preserving local details, providing fine-grained visual information for LLM. See Appendix 11 for algorithm details.
4.3 GlobalCom2 without [CLS] Token
For HR-LVLMs without [CLS] token (e.g., LLaVA-OneVision [21] with SigLIP [50]), we propose an alternative token informativeness measure for GlobalCom2. Specifically, given a sequence of vision tokens $\mathbf{X}\in\mathbb{R}^{N\times d}$ after vision encoding (both thumbnail and crops), we first compute a global mean vector $\mathbf{g}\in\mathbb{R}^{d}$ through global average pooling over all tokens, and then calculate the cosine similarity between each patch token $\mathbf{x}_{i}$ and $\mathbf{g}$:
$c_{i}=\cos(\mathbf{x}_{i},\mathbf{g})=\frac{\mathbf{x}_{i}\cdot\mathbf{g}}{\|\mathbf{x}_{i}\|\|\mathbf{g}\|},$ (7)
The informativeness score $s_{i}=-c_{i}$ is negatively correlated with the similarity, reflecting that tokens with greater difference from $\mathbf{g}$ carry more unique information. Specifically, tokens exhibiting low similarity to $\mathbf{g}$ represent distinctive and irreplaceable visual elements, while highly similar tokens typically correspond to redundant or common patterns. This scoring mechanism serves as an effective alternative to [CLS]-based scoring in Eq. 2-Eq. 5 for models without [CLS] token. It is adopted to evaluate both crop-level information richness and token-level importance in LLaVA-OneVision. We conduct comprehensive quantitative and qualitative analyses in Appendix 8 to explore how to measure token informativeness without [CLS] token.
@@ 5 Experiments
Method VQAv2 GQA VizWiz SQA VQAT POPE MME MMB MMBCN MM-Vet Average
@@
LLaVA-NeXT-7B [29] 81.8 64.2 57.6 70.1 64.9 86.5 1519.0 67.4 60.6 43.9 100.0%
Ratio=75%, Retain up to 2160 Tokens
FastV [8] 81.1 62.5 55.1 69.3 59.7 86.3 1506.3 67.6 59.0 41.7 97.5%
SparseVLM [53] 81.1 62.6 55.2 68.5 60.3 73.2 1507.8 66.1 58.6 41.9 95.7%
FasterVLM [52] 81.1 63.7 56.5 68.4 59.1 87.5 1533.4 67.5 60.2 38.5 97.4%
GlobalCom2 81.3 63.8 56.5 68.7 62.5 87.8 1548.4 68.0 60.6 40.6 98.8%
Ratio=50%, Retain up to 1440 Tokens
FastV [8] 80.7 61.8 54.9 69.1 59.6 85.5 1490.3 67.4 58.5 41.2 96.8%
SparseVLM [53] 80.9 62.0 55.7 68.1 60.0 73.4 1484.9 65.7 58.9 39.9 95.0%
FasterVLM [52] 80.6 63.4 56.4 69.1 58.9 87.7 1533.3 67.4 60.4 39.6 97.7%
GlobalCom2 80.6 63.9 56.5 68.5 62.3 88.1 1552.9 67.6 60.5 40.4 98.6%
@@
FastV [8] 78.9 60.4 54.2 69.8 58.4 83.1 1477.3 65.6 57.0 41.1 95.3%
SparseVLM [53] 78.9 60.9 55.6 67.5 58.1 71.0 1446.1 63.8 57.0 38.0 92.6%
FasterVLM [52] 78.3 61.3 55.4 67.1 58.8 87.2 1454.6 66.0 58.4 37.8 95.1%
GlobalCom2 79.4 61.4 55.7 68.1 60.9 87.6 1493.5 65.9 58.0 40.7 96.6%
Ratio=10%, Retain up to 288 Tokens
FastV [8] 71.9 55.9 53.1 69.3 55.7 71.7 1282.9 61.6 51.9 33.7 87.3%
SparseVLM [53] 71.6 56.1 53.2 68.6 52.0 63.2 1332.2 54.5 50.7 24.7 82.7%
FasterVLM [52] 74.0 56.9 52.6 66.5 56.5 83.6 1359.2 61.6 53.5 35.0 89.8%
GlobalCom2 76.7 57.1 54.6 68.7 58.4 83.8 1365.5 61.8 53.4 36.4 91.5%
Table 1: Comparison with other training-free LVLM token compression methods with LLaVA-NeXT across 10 benchmarks. “Average” shows the mean performance across 10 benchmarks at different retention ratios (75% to 10%), with best results highlighted.
5.1 Experimental Setting
Benchmark Details. We evaluate GlobalCom2 on various multimodal benchmarks including: VQAv2 [14], GQA [17], VizWiz [15], SQA [33], VQAT [39], AI2D [19], MMStar [7], MME [12], MMBench [32], MMBench-CN [32], MM-Vet [48], and SEED-Bench [20].
Implementation Details. We select two representative HR-LVLMs with dynamic tiling: LLaVA-NeXT [29] and LLaVA-OneVision [21] to evaluate the effects of our GlobalCom2. For hyper-parameters, $\tau$ and $\alpha$ are respectively set as 10 and 0.5 for benchmark evaluations. More implementation details are provided in the Appendix 9.
Comparison Details. We compare GlobalCom2 with FastV [8], SparseVLM [53], and FasterVLM [52] at different retention ratios $R$ (75%, 50%, 25%, 10%). While FastV and SparseVLM compress from the default 2880 tokens, FasterVLM and our method first unpad the token sequences - resulting in fewer preserved tokens given the same $R$.
5.2 Main Results
Results on LLaVA-NeXT. Table 1 compares GlobalCom2 with existing methods on 10 benchmarks using LLaVA-NeXT at different retention ratios. (1) GlobalCom2 outperforms baselines across most benchmarks, maintaining >90% of original performance at all retention ratios. Notably, GlobalCom2 even surpasses the original model on several benchmarks (MMB, MME, POPE) even at 50% retention, suggesting effective preservation of semantically crucial tokens for visual understanding. (2) GlobalCom2 excels at low retention ratios (25% and 10%), surpassing FastV and SparseVLM by 4.2% and 8.6% respectively at 10% retention. On VQAT, our method maintains stable performance while others degrade significantly (e.g., SparseVLM drops 12.9%), demonstrating its effectiveness in fine-grained visual understanding. (3) GlobalCom2 demonstrates effective hallucination mitigation on POPE, even surpassing the original model. While text-guided methods (i.e., FastV and SparseVLM) consistently show hallucination issues, vision-semantic approaches (i.e., our GlobalCom2 and FasterVLM) better preserve crucial visual details by avoiding over-reliance on instructions.
Results on LLaVA-OneVision. Figure 5 further presents GlobalCom2’s performance across benchmarks under varying retention ratios $R$ on LLaVA-OneVision. Generally, model performance correlates with $R$, with more aggressive compression leading to degradation. Vision-centric tasks (e.g., GQA, VizWiz, MME, SEED) show substantial drops with reduced visual tokens, while SQA maintains robust performance even with minimal tokens, suggesting language understanding dominates in scientific reasoning. Notably, GlobalCom2 preserves 90.5% performance at $R=10\%$ while consuming only 35.4% of the original GPU memory, all without any training overhead, demonstrating its effective and efficient token compression.
Figure 5: Results of GlobalCom2 on LLaVA-OneVision.
Figure 6: Visualization of token compression. Row 1 shows input images with thumbnail and crops. Rows 2 and 3 present compressed results from FasterVLM and our GlobalCom2 respectively, where grey masks indicate discarded tokens.
@@ 5.3 Ablation Study and Analysis
We conduct ablation studies on GlobalCom2 with $R=25\%$ across five benchmarks (VQAT, POPE, MME, SQA and MM-Vet) covering various vision-language tasks.
Ablation on Adaptive Compression Adjustment for Local Crops. Table 2 compares different compression adjustment strategies: (a) “Uniform”, a uniform strategy with fixed 25% token retention across all crops serving as our baseline, and three adaptive strategies - (b) “$\mathbf{n_{\text{top-{k}}}}$”, which adjusts compression degrees based on the proportion of thumbnail’s top-$k$ ($k=25\%\times N$) most informative tokens present in each crop region, (c) “Softmax (max)”, which applies softmax normalization over the maximum token importance score $s_{j}^{G}$ within each crop’s corresponding thumbnail region, and (d) “Softmax (sum)” (our choice), which computes softmax over the sum of token importance scores $s_{j}^{G}$ in each crop’s corresponding thumbnail region.
As shown in Table 2, our adopted “Softmax (sum)” strategy achieves the best performance. Despite “Uniform”, the other strategies acknowledge varying importance across crops. However, both $\mathbf{n_{\text{top-{k}}}}$ and “Softmax(max)” strategies only consider the strongest visual representation capability within each crop, rather than evaluating the overall importance of the entire crop to the whole image. In contrast, our adopted “Softmax (sum)” strategy dynamically modulates compression intensity according to each crop’s relative importance to the entire image, thereby preserving richer semantic visual information from high-resolution inputs and enabling LLM to better capture fine-grained visual details.
Method VQAT POPE MME SQA MM-Vet Avg.
Upper Bound, 2880 Tokens
Vanilla 64.9 86.5 1519.0 70.1 43.9 100.0%
Ratio=25%, Retain up to 720 Tokens
Uniform 60.1 87.2 1454.6 67.1 37.8 94.2%
$\mathbf{n_{\text{top-{k}}}}$ 59.8 87.3 1471.5 67.4 35.7 94.5%
Softmax (max) 60.3 87.2 1462.6 67.3 38.4 94.7%
Softmax (sum) 60.6 87.4 1473.3 67.6 39.6 95.6%
Table 2: Effects of different compression adjustment strategies.
Ablation on Holistic Token Evaluation for Local Crops. Table 3 compares different token evaluation strategies for token compression. For the $i$-th token in crop $j$, “Global” and “Local” refer to importance scores $\hat{s}_{j,i}^{G}$ and $s_{j,i}^{L}$ respectively. While both global and local token evaluation strategies are effective, each has its limitations. Local-only evaluation excels at fine-grained tasks (VQAT, POPE) but underperforms on general perception benchmarks (MME, SQA) due to missing global context. Conversely, global-only evaluation maintains good general perception but may overlook crucial local details. GlobalCom2 achieves optimal performance by combining both global and local perspectives, benefiting from their complementary strengths.

### v5

@@ 4.1 Adaptive Compression Adjustment
As shown in bottom-left of Figure 5, GlobalCom2 analyzes each local crop’s semantic contribution and applies adaptive token compression accordingly.
For differentiated compression, we compute each crop’s information richness score $s_{j}^{G}$ by accumulating patch-to-[CLS] attention scores in its corresponding global thumbnail region: $s_{j}^{G}=\sum_{i\in\text{crop}_{j}}s^{G}_{i}$. We then normalize scores with $\tilde{s}_{j}=(s_{j}^{G}-\max(s_{j}^{G}))/\tau$ ($\tau=10$) and compute relative importance weight $\sigma_{j}$ via softmax:
@@
where $\epsilon=10^{-8}$ prevents division by zero. The final retention ratio $r_{j}$ for each crop is adjusted from the preset ratio $R(\%)$ based on its global content importance:
@@
where $\sigma_{j}-\frac{1}{n}$ is deviation from average importance, allocating more tokens to important content and fewer to unimportant, for adaptive content-aware compression.
Through this process, GlobalCom2 allocates compression degrees (i.e., $\{r_{1},r_{2},r_{3},r_{4}\}$ in Figure 5) based on each crop’s information richness from the global view.
Method GQA VizWiz SQA MMB POPE VQAT MME $M$-Vet Average
Upper Bound, 2880 Tokens
LLaVA-NeXT-7B 64.2 57.6 70.1 67.4 86.5 64.9 1519.0 43.9 100.0%
Ratio=50%, Retain up to 1440 Tokens
FastV(ECCV24) 61.8 54.9 69.0 67.4 85.5 59.6 1490.3 37.6 95.5%
PDrop(CVPR25) 63.7 57.9 69.2 67.7 87.9 61.6 1499.6 37.5 97.4%
SparseVLM(ICML25) 63.7 57.2 68.3 67.6 87.9 60.5 1507.2 36.8 96.8%
FasterVLM(2024.12) 63.4 56.4 69.1 67.4 87.7 58.9 1533.3 39.6 97.3%
GlobalCom2 63.9 56.5 68.5 67.6 88.1 62.3 1552.9 40.4 98.5%
Ratio=25%, Retain up to 720 Tokens
FastV(ECCV24) 60.4 54.2 68.8 65.6 83.1 58.4 1477.3 35.4 93.4%
PDrop(CVPR25) 60.3 56.8 68.5 65.6 85.5 59.8 1473.7 31.1 93.3%
SparseVLM(ICML25) 59.9 56.0 67.5 65.6 85.0 58.3 1465.9 38.5 94.6%
FasterVLM(2024.12) 61.3 55.4 67.1 66.0 87.2 58.8 1454.6 37.8 94.8%
GlobalCom2 61.5 55.7 68.1 65.9 87.6 60.9 1493.5 40.7 96.7%
Ratio=10%, Retain up to 288 Tokens
FastV(ECCV24) 55.9 53.1 68.1 61.6 71.7 55.7 1282.9 27.2 85.4%
PDrop(CVPR25) 54.5 54.4 67.7 59.0 77.6 54.4 1262.1 24.0 84.3%
SparseVLM(ICML25) 56.3 52.1 68.5 60.0 80.1 53.9 1334.2 26.5 86.1%
PruMerge(ICCV25) 53.6 54.0 66.4 61.3 60.8 50.6 1149.3 25.5 80.6%
FasterVLM(2024.12) 56.9 52.6 66.5 61.6 83.6 56.5 1359.2 35.0 89.9%
GlobalCom2 57.1 54.6 68.7 61.8 83.8 58.4 1365.5 36.4 91.6%
Table 1: Comparisons with LLaVA-NeXT across image understanding benchmarks. VQAT (TextVQA), MME, $M$-Vet are high-resolution benchmarks. “Average” shows mean performance across benchmarks, with best results highlighted.
@@ 4.2 Holistic Token Evaluation
After determining compression degrees for each crop, GlobalCom2 evaluates token importance for preservation (Figure 5). For each crop, attention between patch tokens and [CLS] yields local importance scores $\{s_{j}^{L}\}_{j=1}^{n}$ from final-layer attention. These only capture within-crop importance, missing cross-region elements. Following observation , GlobalCom2 incorporates global thumbnail context by reshaping 1D attention scores ${s}^{G}$ to 2D format and applying bilinear interpolation to match original dimensions, yielding $\{\hat{s}_{j}^{G}\}_{j=1}^{n}$ sub-maps per crop. The holistic score $s_{j,i}$ for the $i$-th token in the $j$-th crop is:
$s_{j,i}=\alpha\hat{s}_{j,i}^{G}+(1-\alpha)s_{j,i}^{L},$ (5)
where we empirically set $\alpha=0.5$ to give equal consideration to both information sources (ablation study of $\alpha$ is in Appendix Figure 10). The final compression:
$\mathbf{X}^{L}_{j}\rightarrow\mathbf{\hat{X}}^{L}_{j}=\text{TopK}(\mathbf{X}^{L}_{j},s_{j},r_{j}\times N).$ (6)
This holistic evaluation identifies globally significant tokens while preserving local details.
Figure 6: Results of GlobalCom2 on LLaVA-OneVision for image understanding. GlobalCom2 achieves 90.5% average performance with only 10% visial tokens.
4.3 Extension to Video Understanding
Given that VideoLLMs process sequential frames with substantial redundancy, analogous to HR-LVLMs with dynamic cropping, we extend GlobalCom2 for efficient VideoLLMs. For video tokens $\mathbf{V}=\{\mathbf{V}_{j}\}_{j=1}^{T}$, we derive global representation $\mathbf{v}^{g}$ via global average pooling. We compute cosine similarity between token $i$ in frame $j$ and $\mathbf{v}^{g}$:
$s_{j,i}^{G}=-\text{sim}(\mathbf{v}_{j,i},\mathbf{v}^{g})=-\frac{\mathbf{v}_{j,i}\cdot\mathbf{v}^{g}}{\|\mathbf{v}_{j,i}\|\|\mathbf{v}^{g}\|},$ (7)
where global score $s_{j,i}^{G}$ is negative cosine similarity, with lower similarity indicating higher distinctiveness. GlobalCom2 calculates information richness $s_{j}^{G}=\sum_{i\in\text{frame}_{j}}s_{j,i}^{G}$ per frame. Following Equations 3-4, this guides compression intensity ($r_{j}$), preserving more tokens in information-dense frames. For local importance, GlobalCom2 generates frame representation $\mathbf{v}_{j}^{f}$ through pooling and computes local scores $s_{j,i}^{L}$ as negative cosine similarity. Combining both perspectives per Equation (5) yields holistic score $s_{j,i}$. Video compression is:
$\mathbf{V}_{j}\rightarrow\mathbf{\hat{V}}_{j}=\text{TopK}(\mathbf{V}_{j},s_{j},r_{j}\times N).$ (8)
Thus, GlobalCom2 enables efficient VideoLLM inference through frame-wise adaptive compression.
@@ 5 Experiments
Method MVBench LongVideoBench MLVU VideoMME Average
Overall Short Medium Long
Upper Bound, 6272 Tokens
LLaVA-OneVision-7B 56.9 56.4 63.0 58.6 70.3 56.6 48.8 100.0%
Ratio=30%, Retain 1882 Tokens
DyCoke(CVPR25) 56.6 54.7 60.3 56.1 67.1 54.6 46.6 96.5%
Ratio=25%, Retain 1568 Tokens
FastV(ECCV24) 55.5 53.3 59.6 55.3 65.0 53.8 47.0 95.0%
PDrop(CVPR25) 55.3 51.3 57.1 55.5 64.7 53.1 48.7 94.2%
SparseVLM(ICML25) 56.4 53.9 60.7 57.3 68.4 55.2 48.1 97.5%
FrameFusion(ICCV25) 56.0 54.8 61.7 57.5 68.2 55.7 48.6 98.1%
GlobalCom2 57.0 55.4 62.6 58.1 69.2 55.8 49.4 99.3%
Ratio=15%, Retain 941 Tokens
FastV(ECCV24) 51.6 48.3 55.0 48.1 51.4 49.4 43.3 85.0%
PDrop(CVPR25) 53.2 47.6 54.7 50.1 58.7 48.7 45.0 87.4%
SparseVLM(ICML25) 52.9 49.7 57.4 53.4 61.0 52.1 47.0 91.2%
GlobalCom2 54.2 52.9 60.7 55.8 66.0 53.3 48.1 95.3%
Table 2: Comparisons with LLaVA-OneVision across video understanding benchmarks. Each benchmark spans different durations: MVBench (16s), LongVideoBench (1-60min), MLVU (3-120min), and VideoMME-S/M/L (1-3/3-30/30-60min).
5.1 Experimental Setting
We evaluate on LLaVA-NeXT (Liu et al. 2024b) and LLaVA-OneVision (Li et al. 2025) for image understanding, LLaVA-OneVision and Qwen2-VL (Wang et al. 2024) for video understanding. We compare with FastV (Chen et al. 2024b), SparseVLM (Zhang et al. 2025), PDrop (Xing et al. 2025), PruMerge (Shang et al. 2025), FasterVLM (Zhang et al. 2024), and VideoLLM methods DyCoke (Tao et al. 2025) and FrameFusion (Fu et al. 2025) at different retention ratios $R$. For fair comparison with multi-stage methods, we use “equivalent token count” reflecting the average percentage of vision tokens retained across all LLM layers.
5.2 Main Results
Method LongVideo. MLVU VideoMME
Overall Short Medium Long
Upper Bound
Qwen2-VL-7B 56.0 60.4 57.6 70.0 54.6 48.2
Ratio=25%
FastV(ECCV24) Out of Memory (OOM)
DyCoke(CVPR25) 50.5 55.2 51.5 62.1 47.1 45.1
GlobalCom2 51.4 55.9 54.3 64.7 50.1 48.2
Table 3: Comparisons on Qwen2-VL. “LongVideo” is LongVideoBench. We conduct on 4 A100 GPUs.
Method SQA POPE VQAT MME $M$-Vet Avg.
@@ 5.2 Main Results
Vanilla 70.1 86.5 64.9 1519.0 43.9 100.0%
@@
Uniform 67.1 87.2 60.1 1454.6 37.8 94.2%
$\mathbf{n_{\text{top-{k}}}}$ 67.4 87.3 59.8 1471.5 35.7 94.5%
Softmax (max) 67.3 87.2 60.3 1462.6 38.4 94.7%
Softmax (sum) 67.6 87.4 60.6 1473.3 39.6 95.6%
Table 4: Effects of different adjustment strategies. “Uniform” performs no compression adjustment, while the other three strategies enable adaptive compression adjustment.
Image Understanding. Table 1 compares GlobalCom2 with existing methods with LLaVA-NeXT, which demonstrates its three key advantages: (i) Superior performance: GlobalCom2 consistently outperforms all baselines, being the only method that maintains >90% of the original performance across all retention ratios. (ii) Extreme compression robustness: While baseline methods suffer significant degradation at $R=10\%$, GlobalCom2 maintains robust performance and achieves the best results across all benchmarks, outperforming the second-best method by 1.7% on average. Notably, FastV and PDrop demonstrate severe deterioration on POPE, exhibiting clear multi-modal hallucination due to positional bias from attention-guided token selection (Figure 7). (iii) High-resolution excellence: GlobalCom2 demonstrates exceptional performance on high-resolution benchmarks (e.g., VQAT, MME, MM-Vet). Our “global-to-local” guided compression preserves both global semantics and local details, outperforming baselines that suffer from over-compression. Figure 6 further validates GlobalCom2’s effectiveness on LLaVA-OneVision, maintaining 90.5% of the original performance at $R=10\%$ while consuming only 35.4% of the original GPU memory. See Appendix for results on the larger HR-LVLM.
Video Understanding. Table 2 presents GlobalCom2’s extension to video understanding with LLaVA-OneVision, demonstrating two key advantages: (i) Superior VideoLLM performance: GlobalCom2 achieves 2.7% average improvement over VideoLLM-specific method DyCoke (Tao et al. 2025) while using fewer vision tokens. (ii) Advantages on long video understanding: At $R=15\%$, GlobalCom2 achieves 3.2 and 3.3 points higher than the second-best method on LongVideoBench and MLVU, demonstrating the effectiveness of our “global-to-local” guided compression design. Table 3 shows GlobalCom2 significantly outperforms DyCoke on Qwen2-VL, confirming its scalability. Additionally, FastV explicitly computes the full attention matrix in the certain layer, causing OOM issues with high token counts and limiting its practical applications.
Method SQA POPE VQAT MME $M$-Vet Avg.
Upper Bound, 2880 Tokens
Vanilla 70.1 86.5 64.9 1519.0 43.9 100.0%
Ratio=25%, Retain up to 720 Tokens
Local only 67.6 87.4 60.6 1473.3 39.6 95.6%
Global only 67.9 86.4 60.2 1488.5 37.8 94.7%
Global and Local 68.1 87.6 60.9 1493.5 40.7 96.7%
Table 5: Effects of different token evaluation metrics. For the $i$-th token in crop $j$, “Global” and “Local” refer to importance scores $\hat{s}_{j,i}^{G}$ and $s_{j,i}^{L}$ in Equation (5), respectively.
@@ 5.3 Ablation Study and Analysis
Figure 7: Visualization of different token compression methods. Gray masks indicate discarded tokens, where other methods exhibit significant over-compression issues, while GlobalCom2 preserves both global important and local detailed information.
Figure 8: Combination with question-aware methods. “+GlobalCom2” indicates the application of our Adaptive Compression Adjustment strategy under $R=10\%$.
Ablation on Adaptive Compression Adjustment. Table 4 compares four settings: (a) “Uniform” baseline with $R=25\%$ across all crops, and three adaptive strategies: (b) $\mathbf{n_{\text{top-{k}}}}$” adjusting compression based on top-$k$ most informative tokens per crop ($k=25\%\times N$), (c) “Softmax (max)” applying softmax over maximum token importance score $s_{j}^{G}$ in each crop’s thumbnail, and (d) “Softmax (sum)” (our choice) computing softmax over sum of token importance scores $s_{j}^{G}$ in each crop’s thumbnail region. All adaptive strategies outperform uniform compression, with “Softmax (sum)” achieving the best performance. While $\mathbf{n_{\text{top-{k}}}}$ and “Softmax (max)” focus on strongest visual features per crop without considering global importance, “Softmax (sum)” adjusts compression based on each crop’s overall importance to the entire image, preserving more semantic information and helping the LLM capture finer visual details.
Ablation on Holistic Token Evaluation. Table 5 compares different token evaluation strategies. While both strategies are effective, each has limitations: Local-only evaluation excels at fine-grained tasks (VQAT, POPE) but underperforms on general perception benchmarks (MME, SQA) due to missing global context. Global-only evaluation maintains general perception but may overlook crucial local details. GlobalCom2 achieves optimal performance by combining them, leveraging their complementary strengths.
Combination with Question-aware Methods. Figure 8 explores combining GlobalCom2 with question-aware methods FastV and SparseVLM, enabling joint consideration of textual relevance and visual importance. Using our Adaptive Compression Adjustment strategy, we assign optimal compression intensities (i.e., $r_{j}$ for the $j$-th crop) per crop based on its visual information richness within the global context before applying FastV/SparseVLM’s token evaluation metrics. Under extreme compression ($R=10\%$), incorporating GlobalCom2 yields average improvements of 5.3% and 5.2% for FastV and SparseVLM. Notably, GlobalCom2 significantly boosts performance on POPE, with improvements of 8.2 for FastV and 4.5 for SparseVLM, confirming our strategy’s effectiveness in mitigating positional bias by preventing over-compression of important regions.